# Mixed Strategy Nash Equilibrium for Crowd Navigation

### v4

@@ III-A Notations and definitions
While we will fully introduce the game theory formula in Section IV, we introduce the concept of strategy here. A strategy describes an agent’s decision-making process. In the context of navigation, we assume agents make decisions regarding trajectories, which outline where they intend to travel over a given time period $[0,T]$.
@@
III-B Iterative Bayesian update with two agents
We describe the iterative Bayesian update scheme that finds cooperative mixed strategies between two navigation agents. Bayesian update requires two key components: a prior belief and a conditional likelihood function—the latter is often interpreted as a measurement model. Here, the prior belief is each agent’s nominal mixed strategy $p^{\prime}(s)$. By assuming each agent follows the expected utility hypothesis [70], we propose the following conditional likelihood function to reflect collision avoidance behavior.
@@ III-B Iterative Bayesian update with two agents
$z(s|p_{j})$ $\propto\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}.$ (6)
@@
III-C Iterative Bayesian update with multiple agents
@@
$z(s|p_{/i})$ $\propto\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}.$ (9)
$\mathbb{E}_{p_{/i}}[r](s)$ $=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{j}}[r](s).$ (10)
@@
The complete iterative Bayesian update algorithm for multi-agent crowd navigation is described in Algorithm 2. Figure 2(c) shows the evolution of multi-agent mixed strategies in a crossing scenario. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE).
@@
III-D Iterative Bayesian update for model predictive crowd navigation
@@
Computing the analytical mixed strategies following the Bayesian update in Algorithm 2 is intractable in practice, as the mixed strategies are represented as continuous probability density functions in the analytical formula. This is a common issue among Bayesian filter methods, as computing the analytical Bayesian posterior is, in general, intractable. Importance sampling [71] is one of the most widely used solutions to approximate the posterior from a Bayesian update efficiently. It represents the prior belief using samples drawn from the prior belief, then it computes the posterior belief by updating the sample weights based on the measurement model. With the sample weights updated, the samples can be updated further through resampling methods, such as rejection sampling [72], for the next iteration. We apply importance sampling to approximate Algorithm 2 in practice. The mixed strategies are represented as samples and weights are computed based on the Bayesian update step in Algorithm 2. Similar sampling-based representation is also used in stochastic optimal control, such as path integral control [73].
@@
One of the requirements for the sampling-based Bayesian update is that we can draw samples from the prior belief, also often known as the proposal distribution, which in Algorithm 2 is the nominal mixed strategy of each agent. We model the nominal mixed strategy for each agent using Gaussian processes [74]. A Gaussian process is an infinite-dimensional normal distribution over continuous functions. In our case, these continuous functions are the trajectories, as we model a trajectory as a continuous mapping from time to the agent state. A Gaussian process is characterized by a mean function and a covariance kernel function. We use a constant velocity model as the mean function for each pedestrian, assuming the pedestrian will keep the current velocity during the robot’s planning horizon [^2] In [28], the constant velocity model is shown to be competitive against state-of-the-art learning-based models for trajectory prediction [28].. A high-level meta-planner generates the mean function for the robot as a trajectory from the robot’s current location toward the navigation goal. With the presence of static obstacles, the mean function trajectory should also avoid the obstacles. Such trajectory planning problems have been well-studied and can be solved using methods such as RRT or trajectory optimization. The parameters of the covariance kernel function can be learned from existing datasets through standard inference techniques such as maximum likelihood estimation [6] or hand-tuned if the number of parameters is small (e.g., with radial basis kernels). Once a Gaussian process is specified, following the same step in Gaussian process regression (Eq 2.24 in [74]), we can condition the specified Gaussian process on discrete time steps. This conditioning happens offline and converts the infinite-dimensional Gaussian process to a finite-dimensional normal distribution with each dimension representing a time step. It also allows us to specify the marginal uncertainty at specific time steps. The converted multivariate normal distribution is an agent’s nominal mixed strategy in practice, from which trajectory samples can be drawn efficiently during runtime, as the sampling of trajectories is now equivalent to sampling from a normal distribution. Figure 3 illustrates the steps of specifying a nominal mixed strategy with a Gaussian process. Lastly, Gaussian processes do not explicitly model agents’ dynamics. Instead, the Gaussian process kernel preserves the geometrical smoothness of the trajectory samples. Similar Gaussian process-based trajectory distribution representations have been verified in other works for crowd navigation [6] and motion planning [75].
@@
Even though we use a Gaussian process-based model for the mixed strategy, it is not the only choice. For pedestrians, the nominal mixed strategy can be the predicted trajectory samples drawn from neural network-based trajectory prediction frameworks, such as Trajectron [26] or TrajNet [76]. For the robot, instead of directly drawing samples in the trajectory space, we can also draw samples in the space of control signals by randomly perturbing a nominal control signal, similar to the widely used model predictive path integral (MPPI) control framework [73, 77]. We choose Gaussian processes for computation and sampling efficiency. Inference with Gaussian processes, which computes both the mean function and the covariance kernel from existing data, is known to be computationally expensive. However, we avoid performing Gaussian process inference by specifying the mean functions and the kernel parameters in advance. Therefore, drawing trajectory samples is equivalent to drawing samples from a multivariate normal distribution, which is affordable for real-time computation and faster than inference from neural networks.
@@
We analyze the computational time complexity of a single iteration within the sampling-based algorithm. Given $T$ time steps as the planning horizon, $N$ agents, and $M$ samples for each agent, the complexity of computing the weights for one agent’s mixed strategy is $O(TMN)$. For each iteration of the Bayesian update scheme, the weights need to be computed for all the agents, which leads to an overall complexity of $O(TMN^{2})$ for one iteration. Note that, given $N$ agents, $O(N^{2})$ is the minimal complexity required for reasoning over all possible two-agent interaction pairs. The computation complexity of our algorithm with respect to the number of agents is lower than the state-of-the-art dynamic game solver ALGAMES [60], which has a computation complexity of $O(TN^{3})$.
IV Mixed strategy Nash equilibrium for crowd navigation
@@
IV-A Preliminaries on game theory
@@
Based on the above definition, it is clear that each agent cannot optimize their individual objective function in isolation from other agents. Thus, the conventional optimality criteria for single-agent optimal decision-making no longer apply. Instead, Nash equilibrium [10, 11] is proposed to describe an equilibrium state where each agent’s strategy is optimal with respect to all other agents’ current strategies.
@@
$J_{i}(p_{1},\cdots,p_{N})$ $=\mathbb{E}_{p_{i},p_{/i}}[r]+D(p_{i}\|p^{\prime}_{i})$ (18)
$\mathbb{E}_{p_{i},p_{/i}}[r]$ $=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{i},p_{j}}[r]$ (19)
@@
IV-B Mixed strategy Nash equilibrium for crowd navigation
@@
$\forall\epsilon>0,\exists K\in\mathbb{N},$
@@
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier. Based on the Lagrangian, the necessary condition for $p^{*}(s)$ to be an extreme is (Theorem 1, Page 43 [78]):
@@
By substituting (26) to the equation constraint (23), we have:
@@
The exponential-based weight update scheme, as well as its mathematical optimality, are not unique to crowd navigation and are not restricted to the formula in this paper. For example, in [79] the same weight update scheme is discussed in the context of online learning.
@@
IV-C Connections to stochastic control and stochastic game-theoretic methods
@@
The derivation of Theorem 2 and Theorem 3 is equivalent to the derivation of the information-theoretic duality between free energy and relative entropy (KL-divergence) in stochastic control [80]. In particular, if we specify the mixed strategy as a stochastic control process and specify the nominal mixed strategy as the path distribution of an uncontrolled system, then the individual objective function (18) is the upper bound of the agent’s free energy. In this case, Theorem 2 and Theorem 3 prove that each BRNE iteration globally minimizes the upper bound of each agent’s free energy, connecting our theoretical results to other results in [80].
@@
Our BRNE algorithm is based on two fundamental numerical techniques: iterative best response (IBR) and importance sampling. IBR is a widely used technique to efficiently approximate the Nash equilibrium of a game. Importance sampling is also widely used in stochastic control, particularly path integral control frameworks [73], to find the optimal control policy. In [61], a model predictive path integral control framework is combined with IBR (BR-MPPI) to find the optimal stochastic control policy for a vehicle interacting with other vehicles. While our BRNE framework and BR-MPPI share similar numerical techniques, there are two fundamental differences: (1) BR-MPPI is a general-purpose game-theoretic planner and BRNE is a game-theoretic crowd navigation planner; (2) Benefitting from its narrow scope, BRNE guarantees the convergence to the global Nash equilibrium and has sufficient computation efficiency to interact with a relatively large number of agents in crowds, while BR-MPPI does not have these advantages when applied to crowd navigation.
@@
The stochastic formulation of BRNE is motivated by the uncertain nature of human behavior—one main advantage of the stochastic formula is that BRNE could capture the multimodality in an agent’s strategy (see Figure 2 for examples). Similarly, a stochastic game-theoretic planner (MPOGames) is proposed in [63], which explicitly computes multi-modal control policies in multi-agent interaction. The key difference between BRNE and MPOGames is the assumed origin of the multimodality in agent intents. The multimodality in MPOGames comes from the assumption that non-ego agents’ objectives are only partially observable to the ego agent. As a result, the ego agent maintains a belief over possible non-ego agent objectives using the Bayesian update, which creates multimodality. On the other hand, in BRNE, all agents’ objectives are assumed to be shared among all agents. The multimodality in BRNE comes from the global Nash equilibrium of a mixed strategy game recovered by BRNE, where the recovered Nash equilibrium could contain multimodal mixed strategies. Essentially, MPOGames and BRNE are making orthogonal contributions: MPOGames paves the way for extending BRNE to scenarios with incomplete information regarding agent objectives, and BRNE offers a potential solution to extend MPOGames beyond Gaussian-based policies and local approximation of Nash equilibrium.
V Evaluation
@@
TABLE I: Safety metrics of multi-agent navigation experiments.
@@
TABLE II: Path efficiency of multi-agent navigation experiments.
@@
V-A Overview of evaluation
We evaluate our BRNE navigation method, as well as other non-learning and learning-based crowd navigation methods, in two categories of tasks: (1) multi-agent navigation (Section V-B); (2) crowd navigation (Section V-C, Section V-D). In the multi-agent navigation experiment, the tested methods are asked to plan paths for all the agents, given a set of start and goal locations. The multi-agent navigation experiment represents the most ideal situation of crowd navigation, where all agents’ behaviors follow the behavioral model assumed by the planning algorithm. The results of the multi-agent navigation experiment do not necessarily reflect an algorithm’s crowd navigation performance with humans, but serve as an upper bound of the algorithms’ ability to infer safe and efficient joint paths—if a method cannot infer safe and efficient joint paths even when all the agents follow the assumed behavioral model, then the method is unlikely to generate safe and efficient navigation actions with real humans who do not necessarily follow the assumed behavioral model. To evaluate the crowd navigation performance of the tested methods, especially when pedestrians do not follow the behavior assumed by the planning algorithm, we conduct extensive crowd navigation benchmark experiments with two different kinds of experiment designs. The first crowd navigation benchmark experiment (Section V-C) places all the agents in a tight space where collision avoidance is necessary for all the agents. The pedestrian behavior is simulated using a decentralized collision avoidance algorithm ORCA [8]. The purpose of this evaluation is to stress test the robot’s ability to infer and adapt to cooperative but unknown joint navigation strategies from other agents. The second crowd navigation benchmark experiment (Section V-D) aims to benchmark the algorithms’ performance in realistic crowd navigation scenarios. We evaluate the performance using a state-of-the-art crowd navigation benchmark [81], which consists of real-world environmental geometry and real human navigation trajectories collected from several datasets. The robot has to navigate in unstructured spaces while interacting with human pedestrians. For each experiment, we evaluate the algorithm performance from two categories of metrics: safety and navigation efficiency. Details of the specific metrics are provided in the corresponding subsections. Lastly, we implement the algorithm on an untethered quadruped robot with fully onboard perception and computation to demonstrate our algorithm’s practicality with real humans. Videos of the benchmark experiments and hardware demonstration are included in the project website: https://sites.google.com/view/brne-crowdnav. We will also release the code and parameters of our implementation.
V-B Multi-agent navigation evaluation

### v6

@@ 3.1 Notations and definitions
While we will fully introduce the game theory formula in Section 4, we introduce the concept of strategy here. A strategy describes an agent’s decision-making process. In the context of navigation, we assume agents make decisions regarding trajectories, which outline where they intend to travel over a given time period $[0,T]$.
@@
3.2 Iterative Bayesian update with two agents
We describe the iterative Bayesian update scheme that finds cooperative mixed strategies between two navigation agents. Bayesian update requires two key components: a prior belief and a conditional likelihood function—the latter is often interpreted as a measurement model. Here, the prior belief is each agent’s nominal mixed strategy $p^{\prime}(s)$. By assuming each agent follows the expected utility hypothesis (Von Neumann and Morgenstern, 1947), we propose the following conditional likelihood function to reflect collision avoidance behavior.
@@ 3.2 Iterative Bayesian update with two agents
$\color[rgb]{0,0,0}z(s|p_{j})$ $\color[rgb]{0,0,0}\propto\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}.$ (6)
@@
3.3 Iterative Bayesian update with multiple agents
@@
$\color[rgb]{0,0,0}z(s|p_{/i})$ $\color[rgb]{0,0,0}\propto\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}.$ (9)
$\color[rgb]{0,0,0}\mathbb{E}_{p_{/i}}[r](s)$ $\color[rgb]{0,0,0}=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{j}}[r](s).$ (10)
@@
The complete iterative Bayesian update algorithm for multi-agent crowd navigation is described in Algorithm 2. Figure 2(c) shows the evolution of multi-agent mixed strategies in a crossing scenario. We name our algorithm Bayesian Recursive Nash Equilibrium (BRNE).
@@
3.4 Iterative Bayesian update for model predictive crowd navigation
@@
Computing the analytical mixed strategies following the Bayesian update in Algorithm 2 is intractable in practice, as the mixed strategies are represented as continuous probability density functions in the analytical formula. This is a common issue among Bayesian filter methods, as computing the analytical Bayesian posterior is, in general, intractable. Importance sampling (Thrun et al., 2005) is one of the most widely used solutions to approximate the posterior from a Bayesian update efficiently. It represents the prior belief using samples drawn from the prior belief, then it computes the posterior belief by updating the sample weights based on the measurement model. With the sample weights updated, the samples can be updated further through resampling methods, such as rejection sampling (Casella et al., 2004), for the next iteration. We apply importance sampling to approximate Algorithm 2 in practice. The mixed strategies are represented as samples and weights are computed based on the Bayesian update step in Algorithm 2. Similar sampling-based representation is also used in stochastic optimal control, such as path integral control (Theodorou et al., 2010).
@@
One of the requirements for the sampling-based Bayesian update is that we can draw samples from the prior belief, also often known as the proposal distribution, which in Algorithm 2 is the nominal mixed strategy of each agent. We model the nominal mixed strategy for each agent using Gaussian processes (Rasmussen and Williams, 2006). A Gaussian process is an infinite-dimensional normal distribution over continuous functions. In our case, these continuous functions are the trajectories, as we model a trajectory as a continuous mapping from time to the agent state. A Gaussian process is characterized by a mean function and a covariance kernel function. We use a constant velocity model as the mean function for each pedestrian, assuming the pedestrian will keep the current velocity during the robot’s planning horizon [^2] In Schöller et al. (2020), the constant velocity model is shown to be competitive against state-of-the-art learning-based models for trajectory prediction.. A high-level meta-planner generates the mean function for the robot as a trajectory from the robot’s current location toward the navigation goal. With the presence of static obstacles, the mean function trajectory should also avoid the obstacles. Such trajectory planning problems have been well-studied and can be solved using methods such as RRT or trajectory optimization. The parameters of the covariance kernel function can be learned from existing datasets through standard inference techniques such as maximum likelihood estimation (Trautman et al., 2015) or hand-tuned if the number of parameters is small (e.g., with radial basis kernels). Once a Gaussian process is specified, following the same step in Gaussian process regression (Eq 2.24 in Rasmussen and Williams (2006)), we can condition the specified Gaussian process on discrete time steps. This conditioning happens offline and converts the infinite-dimensional Gaussian process to a finite-dimensional normal distribution with each dimension representing a time step. It also allows us to specify the marginal uncertainty at specific time steps. The converted multivariate normal distribution is an agent’s nominal mixed strategy in practice, from which trajectory samples can be drawn efficiently during runtime, as the sampling of trajectories is now equivalent to sampling from a normal distribution. Figure 3 illustrates the steps of specifying a nominal mixed strategy with a Gaussian process. Lastly, Gaussian processes do not explicitly model agents’ dynamics. Instead, the Gaussian process kernel preserves the geometrical smoothness of the trajectory samples. Similar Gaussian process-based trajectory distribution representations have been verified in other works for crowd navigation (Trautman et al., 2015) and motion planning (Mukadam et al., 2018).
@@
Even though we use a Gaussian process-based model for the mixed strategy, it is not the only choice. For pedestrians, the nominal mixed strategy can be the predicted trajectory samples drawn from neural network-based trajectory prediction frameworks, such as Trajectron (Salzmann et al., 2020) or TrajNet (Kothari et al., 2022). For the robot, instead of directly drawing samples in the trajectory space, we can also draw samples in the space of control signals by randomly perturbing a nominal control signal, similar to the widely used model predictive path integral (MPPI) control framework (Theodorou et al., 2010; Williams et al., 2016). We choose Gaussian processes for computation and sampling efficiency. Inference with Gaussian processes, which computes both the mean function and the covariance kernel from existing data, is known to be computationally expensive. However, we avoid performing Gaussian process inference by specifying the mean functions and the kernel parameters in advance. Therefore, drawing trajectory samples is equivalent to drawing samples from a multivariate normal distribution, which is affordable for real-time computation and faster than inference from neural networks.
@@
We analyze the computational time complexity of a single iteration within the sampling-based algorithm. Given $T$ time steps as the planning horizon, $N$ agents, and $M$ samples for each agent, the complexity of computing the weights for one agent’s mixed strategy is $O(TMN)$. For each iteration of the Bayesian update scheme, the weights need to be computed for all the agents, which leads to an overall complexity of $O(TMN^{2})$ for one iteration. Note that, given $N$ agents, $O(N^{2})$ is the minimal complexity required for reasoning over all possible two-agent interaction pairs. The computation complexity of our algorithm with respect to the number of agents is lower than the state-of-the-art dynamic game solver ALGAMES (Le Cleac’h et al., 2022), which has a computation complexity of $O(TN^{3})$.
4 Mixed strategy Nash equilibrium for crowd navigation
@@
4.1 Preliminaries on game theory
@@
Based on the above definition, it is clear that each agent cannot optimize their individual objective function in isolation from other agents. Thus, the conventional optimality criteria for single-agent optimal decision-making no longer apply. Instead, Nash equilibrium (Nash, 1950, 1951) is proposed to describe an equilibrium state where each agent’s strategy is optimal with respect to all other agents’ current strategies.
@@
$\color[rgb]{0,0,0}J_{i}(p_{1},\cdots,p_{N})$ $\color[rgb]{0,0,0}=\mathbb{E}_{p_{i},p_{/i}}[r]+D(p_{i}\|p^{\prime}_{i})$ (18)
$\color[rgb]{0,0,0}\mathbb{E}_{p_{i},p_{/i}}[r]$ $\color[rgb]{0,0,0}=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{i},p_{j}}[r]$ (19)
@@
4.2 Mixed strategy Nash equilibrium for crowd navigation
@@
$\color[rgb]{0,0,0}\forall\epsilon>0,\exists K\in\mathbb{N},$
@@
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier. Based on the Lagrangian, the necessary condition for $p^{*}(s)$ to be an extreme is (Theorem 1, Page 43 (Gelfand et al., 2000)):
@@
By substituting (26) to the equality constraint (23), we have:
@@
The exponential-based weight update scheme, as well as its mathematical optimality, are not unique to crowd navigation and are not restricted to the formula in this paper. For example, in Hoeven et al. (2018) the same weight update scheme is discussed in the context of online learning.
@@
4.3 Connections to stochastic control and stochastic game-theoretic methods
@@
The derivation of Theorem 2 and Theorem 3 is equivalent to the derivation of the information-theoretic duality between free energy and relative entropy (KL-divergence) in stochastic control (Theodorou and Todorov, 2012). In particular, if we specify the mixed strategy as a stochastic control process and specify the nominal mixed strategy as the path distribution of an uncontrolled system, then the individual objective function (18) is the upper bound of the agent’s free energy. In this case, Theorem 2 and Theorem 3 prove that each BRNE iteration globally minimizes the upper bound of each agent’s free energy, connecting our theoretical results to other results in Theodorou and Todorov (2012).
@@
Our BRNE algorithm is based on two fundamental numerical techniques: iterative best response (IBR) and importance sampling. IBR is a widely used technique to efficiently approximate the Nash equilibrium of a game. Importance sampling is also widely used in stochastic control, particularly path integral control frameworks (Theodorou et al., 2010), to find the optimal control policy. In Williams et al. (2018), a model predictive path integral control framework is combined with IBR (BR-MPPI) to find the optimal stochastic control policy for a vehicle interacting with other vehicles. While our BRNE framework and BR-MPPI share similar numerical techniques, there are two fundamental differences: (1) BR-MPPI is a general-purpose game-theoretic planner and BRNE is a game-theoretic crowd navigation planner; (2) Benefitting from its narrow scope, BRNE guarantees the convergence to the global Nash equilibrium and has sufficient computation efficiency to interact with a relatively large number of agents in crowds, while BR-MPPI does not have these advantages when applied to crowd navigation.
@@
The stochastic formulation of BRNE is motivated by the uncertain nature of human behavior—one main advantage of the stochastic formula is that BRNE could capture the multimodality in an agent’s strategy (see Figure 2 for examples). Similarly, a stochastic game-theoretic planner (MPOGames) is proposed in So et al. (2023), which explicitly computes multi-modal control policies in multi-agent interaction. The key difference between BRNE and MPOGames is the assumed origin of the multimodality in agent intents. The multimodality in MPOGames comes from the assumption that non-ego agents’ objectives are only partially observable to the ego agent. As a result, the ego agent maintains a belief over possible non-ego agent objectives using the Bayesian update, which creates multimodality. On the other hand, in BRNE, all agents’ objectives are assumed to be shared among all agents. The multimodality in BRNE comes from the global Nash equilibrium of a mixed strategy game recovered by BRNE, where the recovered Nash equilibrium could contain multimodal mixed strategies. Essentially, MPOGames and BRNE are making orthogonal contributions: MPOGames paves the way for extending BRNE to scenarios with incomplete information regarding agent objectives, and BRNE offers a potential solution to extend MPOGames beyond Gaussian-based policies and local approximation of Nash equilibrium.
5 Evaluation
@@
Table 1: Safety metrics of multi-agent navigation experiments.
@@
Table 2: Path efficiency of multi-agent navigation experiments.
@@
5.1 Overview of evaluation
We evaluate our BRNE navigation method, as well as other non-learning and learning-based crowd navigation methods, in two categories of tasks: (1) multi-agent navigation (Section 5.2); (2) crowd navigation (Section 5.3, Section 5.4). In the multi-agent navigation experiment, the tested methods are asked to plan paths for all the agents, given a set of start and goal locations. The multi-agent navigation experiment represents the most ideal situation of crowd navigation, where all agents’ behaviors follow the behavioral model assumed by the planning algorithm. The results of the multi-agent navigation experiment do not necessarily reflect an algorithm’s crowd navigation performance with humans, but serve as an upper bound of the algorithms’ ability to infer safe and efficient joint paths—if a method cannot infer safe and efficient joint paths even when all the agents follow the assumed behavioral model, then the method is unlikely to generate safe and efficient navigation actions with real humans who do not necessarily follow the assumed behavioral model. To evaluate the crowd navigation performance of the tested methods, especially when pedestrians do not follow the behavior assumed by the planning algorithm, we conduct extensive crowd navigation benchmark experiments with two different kinds of experiment designs. The first crowd navigation benchmark experiment (Section 5.3) places all the agents in a tight space where collision avoidance is necessary for all the agents. The pedestrian behavior is simulated using a decentralized collision avoidance algorithm ORCA (van den Berg et al., 2011). The purpose of this evaluation is to stress test the robot’s ability to infer and adapt to cooperative but unknown joint navigation strategies from other agents. The second crowd navigation benchmark experiment (Section 5.4) aims to benchmark the algorithms’ performance in realistic crowd navigation scenarios. We evaluate the performance using a state-of-the-art crowd navigation benchmark (Biswas et al., 2022), which consists of real-world environmental geometry and real human navigation trajectories collected from several datasets. The robot has to navigate in unstructured spaces while interacting with human pedestrians. For each experiment, we evaluate the algorithm performance from two categories of metrics: safety and navigation efficiency. Details of the specific metrics are provided in the corresponding subsections. Lastly, we implement the algorithm on an untethered quadruped robot with fully onboard perception and computation to demonstrate our algorithm’s practicality with real humans. The implementation of the algorithm, including the parameters and tutorials, will be released on https://sites.google.com/view/brne-crowdnav.
5.2 Multi-agent navigation evaluation